OpenStudy (anonymous):

solve the problem and show the steps 4w-4=w^2

OpenStudy (anonymous):

w^2-4w+4=0 (w-2)^2=0 w-2=0 w=2

OpenStudy (anonymous):

wouldnt it be -w^2?

OpenStudy (anonymous):

no i took left side terms to right side

OpenStudy (anonymous):

oooo ok i see it. thanks!

OpenStudy (anonymous):

you're welcome :)